The power produces by a windmill varies directly with the square of its diameter and the cube of the wind speed. A fan with a diameter of 2.5 meters produces 560 watts of power if the wind speed is 4 m/s. How much power would a similar fan produce if it had a diameter of 5 meters and the wind speed was 5 m/s?
Answer by Alan3354(69443)

The power produces by a windmill varies directly with the square of its diameter and the cube of the wind speed. A fan with a diameter of 2.5 meters produces 560 watts of power if the wind speed is 4 m/s. How much power would a similar fan produce if it had a diameter of 5 meters and the wind speed was 5 m/s?
-------------
P = 560*(5/2.5)^2*(5/4)^3 = 560*4*125/64
P = 4375 watts